# Nowcast3D: Reliable precipitation nowcasting via gray-box learning

### v3

@@ Methods
where the full velocity field is reconstructed from a scalar potential $\varphi$ and a vector potential $\boldsymbol{\psi}$ via $\boldsymbol{v}_{1:T}=\nabla\varphi_{1:T}+\nabla\times\boldsymbol{\psi}_{1:T}.$ The backbone parameters are shared across all outputs, whereas the three decoders have separate parameters and specialise in inferring the dynamics of motion, diffusion and source terms, respectively. Additional architectural details are provided in Supplementary Figure 2.
@@
Given the inferred fields $(\boldsymbol{v}_{1:T},\boldsymbol{\kappa}_{1:T},s_{1:T})$, we advance the radar field $R_{t}$ forward in time by numerically integrating Equation (2). We use an operator-splitting scheme that decomposes each time step into successive advection, diffusion and source updates. The advection operator
@@
is implemented with a semi-Lagrangian method, which remains stable for large Courant numbers and is well suited to long lead times. To implement anisotropic diffusion, we exploit the link between diffusion and random motion [einstein1905molekularkinetischen, chapman1990mathematical, zhang2025monte] and define the diffusion operator as
$\mathcal{D}_{\Delta t}[R_{t}](x)=\mathbb{E}_{\eta\sim\mathcal{N}\bigl(0,\,2\boldsymbol{\kappa}_{t}\Delta t\bigr)}\left[R_{t}(x+\eta)\right]$ (5)
@@
Composing these operators yields the one-step update
@@
where each stochastic displacement $\eta^{(m)}_{t}(x)$ is drawn independently from $\mathcal{N}\big(0,\,2\boldsymbol{\kappa}_{t}(x)\Delta t\big)$.Throughout both training and inference, we fix the number of diffusion samples to $M=8$.
@@
Because the physics-informed module already advances the full three-dimensional reflectivity field through explicit dynamical operators, the role of $\boldsymbol{\Phi}$ is to provide a stochastic refinement around this baseline. To reduce training complexity and computational cost, we implement this refinement in two-dimensional space, operating on full time sequences of radar cross-sections rather than on complete 3D volumes. The same trained network can subsequently be applied either to column-maximum reflectivity sequences to generate 2D probabilistic forecasts, or independently to sequences from multiple height levels, whose outputs can be stacked to reconstruct a three-dimensional ensemble of reflectivity volumes. Further construction and sampling details are provided in Supplementary Note 2.2.
@@
Operationally, the probabilistic refinement module implements the sampling step $R^{(k)}_{1:T}=\mathrm{Sample}_{\boldsymbol{\Phi}}(\epsilon^{(k)},c)$ used to construct the ensemble forecast. For each noise realisation $\epsilon^{(k)}$, the two diffusion branches produce a pair of correlated outputs: a full-field structure $S_{\mathrm{struct}}^{(k)}$ and a residual field $S_{\mathrm{res}}^{(k)}$ defined relative to the deterministic forecast. This pair of outputs supports three types of ensemble members constructed from each sample:
@@
where $\alpha\in[0,1]$ is a mixing coefficient. Generating $K$ independent noise realisations yields $K$ pairs $\{(S_{\mathrm{struct}}^{(k)},S_{\mathrm{res}}^{(k)})\}_{k=1}^{K}$, from which we form the ensemble
@@
where $S^{(i)}_{1:T}(k)$ denotes the $i$-th construction in (8) for sample $k$. In the hybrid formulation, all these members are collectively denoted by $R^{\mathrm{ens}}_{1:T}$ and are used for probabilistic evaluation. The ensemble forecasts shown in the figures correspond to the mean over all ensemble members (see Supplementary Note 2.3 for additional implementation details).
@@
Training proceeds in two stages that mirror the hybrid formulation described above. In the first stage, we train the physics-informed module $\boldsymbol{\Theta}$ to produce a deterministic evolution of the radar reflectivity field. In the second stage, we freeze $\boldsymbol{\Theta}$ and train the probabilistic refinement module $\boldsymbol{\Phi}$ to learn a conditional distribution of futures given both the past observations and the deterministic forecast.
@@
In the first stage, we optimise all parameters of the physics-informed module $\boldsymbol{\Theta}$, including the shared backbone and the three task-specific decoders. The objective $\mathcal{L}_{\mathrm{phys}}$ is a composite loss that supervises the intermediate states of the operator sequence as well as the final forecast. At each forecast step $t$, we penalise discrepancies between (i) the purely advected field, (ii) the advected–diffused field and (iii) the final source-corrected forecast and the corresponding ground-truth radar volume at time $t+1$. The total loss is a weighted sum of $\ell_{1}$ norms accumulated over the forecast horizon,
@@
Here $R_{t}$ and $R_{t+1}$ denote the model forecasts at times $t$ and $t+1$, and $R_{t+1}^{\text{truth}}$ is the corresponding observed radar volume. The weights $\lambda_{\text{adv}}$ and $\lambda_{\text{diff}}$ balance the contribution of the intermediate losses, encouraging the network to learn physically meaningful advection and diffusion operators in addition to accurate end-of-step forecasts. Further training details are provided in Supplementary Note 3.1.
@@
In the second stage, we train the probabilistic refinement module $\boldsymbol{\Phi}$ while keeping $\boldsymbol{\Theta}$ fixed. Its dual-branch architecture is designed to produce the structured field $S_{\mathrm{struct}}$ and residual field $S_{\mathrm{res}}$ used in the ensemble construction in Equation (8): the structure branch learns a distribution over full future reflectivity fields, whereas the residual branch learns a distribution over corrections to the deterministic forecast.
We train both branches on two-dimensional radar fields rather than on full 3D volumes. For each spatio–temporal window we treat either a single height level or the column-maximum projection as a 2D sample, yielding an augmented training set $R^{\mathrm{2D}}$ with corresponding historical, deterministic and truth sequences $R^{\mathrm{2D}}_{\text{hist}}$, $R^{\mathrm{2D,det}}_{1:T}$ and $R^{\mathrm{2D,truth}}_{1:T}$.
Each branch is trained as a conditional diffusion model using a shared noise schedule and the $v$-parameterisation [salimansprogressive]. For notational simplicity, we write $R_{\text{truth}}$ for $R^{\mathrm{2D,truth}}_{1:T}$ and $r_{\text{truth}}=R_{\text{truth}}-R^{\mathrm{2D,det}}_{1:T}$ for the corresponding residual composites. The two-dimensional conditioning context is $c=\bigl(R^{\mathrm{2D}}_{-T_{0}:0},\,R^{\mathrm{2D,det}}_{1:T}\bigr)$. For each branch, a dedicated context encoder processes this context and produces a hierarchy of multi-scale feature maps that are used as conditioning inputs to the corresponding denoiser. The structure and residual branches are trained with the following diffusion losses,
@@
where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ is Gaussian noise, $\bar{\alpha}_{t}$ is the cumulative product of the noise schedule, and $v_{\theta_{\text{struct}}}$ and $v_{\theta_{\text{res}}}$ denote the predicted $v$-parameters for the structure and residual branches, respectively. Further training details are provided in Supplementary Note 3.2.
@@
We evaluate model performance at two spatial resolutions: 0.04° (coarse) and 0.01° (fine). At both resolutions we use $256\times 256$-pixel input patches. For the 0.04° experiments, this corresponds to a $10.24^{\circ}\times 10.24^{\circ}$ domain that is representative of province-scale regions. For the 0.01° experiments, it corresponds to a $2.56^{\circ}\times 2.56^{\circ}$ domain tailored to high-resolution, city-scale hazards.
@@
We assess forecast quality using three complementary metrics. (i) Event detection and localisation are quantified using neighbourhood CSI at reflectivity thresholds of 20, 30, 40 and 50 dBZ. This metric rewards forecasts that place high-intensity cells close to their observed locations while allowing for small spatial offsets. (ii) Perceptual and structural fidelity are evaluated using the LPIPS metric with a pre-trained AlexNet [krizhevsky2012imagenet] backbone, which correlates well with human visual judgement of image similarity. (iii) Scale-dependent realism is diagnosed using PSD analysis: we compute the two-dimensional power spectrum of each forecast and observation, perform radial averaging to obtain a one-dimensional spectrum as a function of wavelength, and compare spectral slopes and power ratios from convective to mesoscale ranges.
For a fair comparison with two-dimensional baselines, all three metrics are computed on column-maximum reflectivity fields obtained by applying the same vertical projection to the three-dimensional Nowcast3D forecasts and to the input radar volumes used for the 2D models.
@@
To evaluate the physical realism of the velocity field $\boldsymbol{v}$ inferred by Nowcast3D (Equation 3), we compare it with independent observations from wind-profiler radar. This requires a two-step post-processing procedure (Extended Data Figure 5a). First, the model output, expressed as grid-cell displacement per 6-minute interval in degrees, is converted to zonal and meridional velocity components in $\mathrm{m\,s^{-1}}$ using latitude-dependent scaling to account for spherical geometry. Second, because profiler sites do not coincide exactly with model grid points, we perform spatio-temporal matching by defining, for each observation, a horizontal search radius and a vertical tolerance window (Extended Data Figure 5b,c). The corresponding model value is obtained by averaging all grid cells within this three-dimensional cylinder at the matching forecast time.
From the resulting matched pairs, we compute the Pearson correlation coefficient between predicted and observed wind speeds over the forecast horizon (Extended Data Figure 5d), as well as the mean absolute error of wind direction (Extended Data Figure 5e). This analysis provides an independent check on the consistency of the inferred flow fields with observed winds in precipitating regions and complements the reflectivity-based verification described above.
The velocity field derived from the stream-function and potential-function decomposition is detailed in Supplementary Note 6.2.
@@ Acknowledgement
The work is supported by the National Natural Science Foundation of China (No. 62276269, No. 92270118) and the Beijing Natural Science Foundation (No. 1232009). We sincerely thank Professor Xiaoding Yu, for his important contribution to the meteorologist evaluation process. We also extend our heartfelt thanks to the 160 anonymous meteorologists who participated in the meteorologist evaluation.
@@
Extended Data Figure 1: Overview of the Nowcast3D framework. The pipeline consists of three main components: data processing (green), a physics-based prediction network (blue) and a diffusion-based generative refinement network (red).
Extended Data Figure 2: Radar network coverage and low-level imputation examples. a, Spatial distribution of the operational weather radar network over China and surrounding regions, with coloured disks indicating approximate horizontal coverage at 500, 1500, 3000, 4500, 6500 and 10,000 m above sea level. The dashed rectangle marks the analysis domain used in this study. b, Schematic of the three-dimensional radar sampling geometry, illustrating how the beam widens with range and leads to sparse coverage near the surface despite dense sampling aloft. c, Example of low-level reflectivity in the Huanan area at heights of 500 m and 1,000 m before (left) and after (right) imputation. d, Same as c but for a convective event near Maoming, Guangdong, highlighting the recovery of coherent low-level structures beneath well-observed upper-level echoes.
@@
Extended Data Figure 3: 3D Layered Comparison This figure corresponds to the layered prediction map in Figure 2. We show a comparison between our prediction results and the ground truth at a specific vertical height. The 500m observation exhibits circular features due to the limitations of the radar detection range, which appears as an overlap of multiple observation circles at low altitudes.
@@
Extended Data Figure 4: 3D Layered Comparison This figure corresponds to the layered prediction map in Figure 3. We show a comparison between our prediction results and the ground truth at a specific vertical height. The 500m observation exhibits circular features due to the limitations of the radar detection range, which appears as an overlap of multiple observation circles at low altitudes.
@@
Extended Data Figure 5: Wind Field Calculation and Evaluation Methodology. a, Schematic of the calculation process from gridded data to the site-specific wind field. b, Spatial averaging method for evaluation: Since observation sites do not align perfectly with grid points, we define a circular area with a radius of Threshold around each site. The predicted values within this circle are averaged for comparison with the site’s observation. c, Vertical averaging method for evaluation: As observations may not match the exact vertical levels of the predictions, we average all observations within a vertical range of ±Threshold/2 from a specific height for comparison. d, Temporal evolution of the correlation coefficient between predicted and observed wind speeds. e, Temporal evolution of the mean absolute error in wind direction between predicted and observed fields.

### v4

@@ Methods
where the full velocity field is reconstructed from a scalar potential $\varphi$ and a vector potential $\boldsymbol{\psi}$ via $\boldsymbol{v}_{1:T}=\nabla\varphi_{1:T}+\nabla\times\boldsymbol{\psi}_{1:T}.$ The backbone parameters are shared across all outputs, whereas the three decoders have separate parameters and specialize in inferring the dynamics of motion, diffusion and source terms, respectively. Additional architecture details are given in Supplementary Figure 2.
@@
Given the inferred fields $(\boldsymbol{v}_{1:T},\boldsymbol{\kappa}_{1:T},s_{1:T})$, we advance the radar field $R_{t}$ forward in time by numerically integrating Eq. (2). We use an operator-splitting scheme that decomposes each time step into successive advection, diffusion and source updates. The advection operator that reads
@@
is implemented using a semi-Lagrangian method [staniforth1991semi, zhang2023skilful], which remains stable for large Courant numbers and is well suited to long lead times. To implement anisotropic diffusion, we exploit the link between diffusion and random motion [einstein1905molekularkinetischen, chapman1990mathematical, zhang2025monte] and define the diffusion operator as
$\mathcal{D}_{\Delta t}[R_{t}](x)=\mathbb{E}_{\eta\sim\mathcal{N}\bigl(0,\,2\boldsymbol{\kappa}_{t}\Delta t\bigr)}\left[R_{t}(x+\eta)\right],$ (5)
@@
Composing these operators yields the one-step update expressed as
@@
where each stochastic displacement $\eta^{(m)}_{t}(x)$ is drawn independently from $\mathcal{N}\big(0,\,2\boldsymbol{\kappa}_{t}(x)\Delta t\big)$. Throughout both training and inference, we fix empirically the number of diffusion samples to $M=8$.
@@
Because the physics-informed learning module already advances the full 3D reflectivity field through explicit dynamical operators, the role of $\boldsymbol{\Phi}$ is to provide a stochastic refinement around this baseline state. To reduce training complexity and computational cost, we implement this refinement in the 2D space, operating on full time sequences of radar cross-sections rather than on complete 3D volumes. The same trained network can subsequently be applied either to column-maximum reflectivity sequences to generate 2D probabilistic forecasts, or independently to sequences from multiple height levels, whose outputs can be stacked to reconstruct a 3D ensemble of reflectivity volumes. Further construction and sampling details are provided in Supplementary Note 2.2.
@@
Operationally, the probabilistic refinement module implements the sampling step $R^{(k)}_{1:T}=\mathrm{Sample}_{\boldsymbol{\Phi}}(\epsilon^{(k)},c)$ used to construct the ensemble forecast. For each noise realisation $\epsilon^{(k)}$, the two diffusion branches produce a pair of correlated outputs: a full-field structure $S_{\mathrm{struct}}^{(k)}$ and a residual field $S_{\mathrm{res}}^{(k)}$ defined relative to the deterministic forecast. This pair of outputs supports three types of ensemble members constructed from each sample, namely,
@@
where $\alpha\in[0,1]$ is a mixing coefficient. Generating $K$ independent noise realisations yields $K$ pairs $\{(S_{\mathrm{struct}}^{(k)},S_{\mathrm{res}}^{(k)})\}_{k=1}^{K}$, from which we form the ensemble:
@@
where $S^{(i)}_{1:T}(k)$ denotes the $i$-th construction in Eq. (8) for sample $k$. In the hybrid formulation, all these members are collectively denoted by $R^{\mathrm{ens}}_{1:T}$ and are used for probabilistic evaluation. The ensemble forecasts shown in the result figures correspond to the mean over all ensemble members (see Supplementary Note 2.3 for additional implementation details).
@@
The model training proceeds in two stages that mirror the hybrid formulation described above. In the first stage, we train the physics-informed module $\boldsymbol{\Theta}$ to produce a deterministic evolution of the radar reflectivity field. In the second stage, we freeze $\boldsymbol{\Theta}$ and train the probabilistic refinement module $\boldsymbol{\Phi}$ to learn a conditional distribution of futures given both the past observations and the deterministic forecast.
@@
In the first stage, we optimize all parameters of the physics-informed module $\boldsymbol{\Theta}$, including the shared backbone and the three task-specific decoders. The objective $\mathcal{L}_{\mathrm{phys}}$ is a composite loss that supervises the intermediate states of the operator sequence as well as the final forecast. At each forecast step $t$, we penalize discrepancies between (i) the purely advected field, (ii) the advected–diffused field and (iii) the final source-corrected forecast and the corresponding ground-truth radar volume at time $t+1$. The total loss is a weighted sum of $\ell_{1}$ norms accumulated over the forecast horizon, namely,
@@
Here, $R_{t}$ and $R_{t+1}$ denote the model forecasts at times $t$ and $t+1$, respectively, and $R_{t+1}^{\text{truth}}$ is the corresponding observed radar volume. The weights $\lambda_{\text{adv}}$ and $\lambda_{\text{diff}}$ balance the contribution of the intermediate losses, encouraging the network to learn physically meaningful advection and diffusion operators in addition to accurate end-of-step forecasts. Further training details are provided in Supplementary Note 3.1.
@@
In the second stage, we train the probabilistic refinement module $\boldsymbol{\Phi}$ while keeping $\boldsymbol{\Theta}$ fixed. Its dual-branch architecture is designed to produce the structured field $S_{\mathrm{struct}}$ and residual field $S_{\mathrm{res}}$ used in the ensemble construction in Eq. (8): the structure branch learns a distribution over full future reflectivity fields, whereas the residual branch learns a distribution over corrections to the deterministic forecast.
We train both branches on 2D radar fields rather than on full 3D radar volumes. For each spatiotemporal window, we treat either a single height level or the column-maximum projection as a 2D sample, yielding an augmented training set $R^{\mathrm{2D}}$ with corresponding historical, deterministic and truth sequences $R^{\mathrm{2D}}_{\text{hist}}$, $R^{\mathrm{2D,det}}_{1:T}$ and $R^{\mathrm{2D,truth}}_{1:T}$.
Each branch is trained as a conditional diffusion model using a shared noise schedule and the $v$-parameterization [salimansprogressive]. For notational simplicity, we write $R_{\text{truth}}$ for $R^{\mathrm{2D,truth}}_{1:T}$ and $r_{\text{truth}}=R_{\text{truth}}-R^{\mathrm{2D,det}}_{1:T}$ for the corresponding residual composites. The 2D conditioning context is given by $c=\bigl(R^{\mathrm{2D}}_{-T_{0}:0},\,R^{\mathrm{2D,det}}_{1:T}\bigr)$. Here, $\tau$ indexes the diffusion step. We sample $\tau\sim\mathrm{Unif}\{0,\dots,N-1\}$ and $\epsilon\sim\mathcal{N}(0,\mathbf{I})$, where $N$ is the total number of diffusion steps. Let $\{\beta_{\tau}\}_{\tau=0}^{N-1}$ be the shared noise schedule, $\alpha_{\tau}=1-\beta_{\tau}$, and $\bar{\alpha}_{\tau}=\prod_{s=0}^{\tau}\alpha_{s}$. For each branch, a dedicated context encoder processes $c$ and produces a hierarchy of multi-scale feature maps used to condition the denoiser; we denote these encoded features by $h_{\text{struct}}=E_{\phi_{\text{struct}}}(c)$ and $h_{\text{res}}=E_{\phi_{\text{res}}}(c)$. The structure and residual branches are trained with the following diffusion losses:
@@
where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ emphasizes the Gaussian noise, $\bar{\alpha}_{\tau}$ is the cumulative product of the noise schedule, and $v_{\theta_{\text{struct}}}$ and $v_{\theta_{\text{res}}}$ denote the predicted $v$-parameters for the structure and residual branches, respectively. Further training details are provided in Supplementary Note 3.2.
@@
We evaluate the model performance at two spatial resolutions: 0.04° (coarse) and 0.01° (fine). At both resolutions, we use $256\times 256$-pixel input patches. For the 0.04° experiments, this corresponds to a $10.24^{\circ}\times 10.24^{\circ}$ domain that is representative of province-scale regions. For the 0.01° experiments, it corresponds to a $2.56^{\circ}\times 2.56^{\circ}$ domain tailored to high-resolution, city-scale hazards.
@@
We assess the forecast quality using three complementary metrics. (i) Event detection and localization are quantified using neighborhood CSI at reflectivity thresholds of 20, 30, 40 and 50 dBZ. This metric rewards forecasts that place high-intensity cells close to their observed locations while allowing for small spatial offsets. (ii) Perceptual and structural fidelity are evaluated using the LPIPS metric with a pre-trained AlexNet [krizhevsky2012imagenet] backbone, which correlates well with human visual judgement of image similarity. (iii) Scale-dependent realism is diagnosed using PSD analysis: we compute the 2D power spectrum of each forecast and observation, perform radial averaging to obtain a 1D spectrum as a function of wavelength, and compare spectral slopes and power ratios from convective to mesoscale ranges.
For a fair comparison with 2D baselines, all three metrics are computed on column-maximum reflectivity fields obtained by applying the same vertical projection to the three-dimensional Nowcast3D forecasts and to the input radar volumes used for the 2D models.
@@
To evaluate the physical realism of the velocity field $\boldsymbol{v}$ inferred by Nowcast3D (see Eq. (3)), we compare it with independent observations from wind-profiler radars. This requires a two-step post-processing procedure (Extended Data Figure 4a). First, the model output, expressed as grid-cell displacement per 6-min interval in degrees, is converted to zonal and meridional velocity components in $\mathrm{m\,s^{-1}}$ using latitude-dependent scaling to account for spherical geometry. Second, because profiler sites do not coincide exactly with our model grid points, we perform spatiotemporal matching by defining, for each observation, a horizontal search radius and a vertical tolerance window (Extended Data Figure 4b,c). The corresponding model value is obtained by averaging all grid cells within this 3D cylinder at the matching forecast time.
From the resulting matched pairs, we compute the Pearson correlation coefficient between predicted and observed wind speeds over the forecast horizon (Extended Data Figure 4d), as well as the mean absolute error of wind direction (Extended Data Figure 4e). This analysis provides an independent check on the consistency of the inferred flow fields with observed winds in precipitating regions and complements the reflectivity-based verification described above. The velocity field derived from the stream-function and potential-function decomposition is detailed in Supplementary Note 6.2.
@@ Acknowledgement
The work is supported by the National Natural Science Foundation of China (62276269, 92270118), the Beijing Natural Science Foundation (1232009), the National Key Research and Development Program of China (2025YFE0217100), and the Strategic Priority Research Program of the Chinese Academy of Sciences (XDB0620103). We sincerely thank Professor Xiaoding Yu, for his important contribution to the meteorologist evaluation process. We also extend our heartfelt thanks to the 160 anonymous meteorologists who participated in the meteorologist evaluation.
@@
Extended Data Figure 1: Overview of the Nowcast3D framework. The pipeline consists of three main components: data processing (green), a physics-based prediction network (blue) and a diffusion-based generative refinement network (red). a, Pipeline of Nowcast3D. Raw 3D radar volumes spanning the period of $T_{\mathrm{in}}=[-T_{0}:0]$ are processed into normalized inputs and an observable mask, while spatiotemporal embeddings are provided as external inputs. Phy-Pred produces a coarse forecast $\mathrm{Radar}^{\mathrm{pred}}$ (1:$T$), which is refined by Diff-Gen with injected noise to generate $\mathrm{Radar}^{\mathrm{gen}}$, followed by ensemble aggregation. b, Module details. Data processing performs mask extraction, masked imputation, and normalization. Phy-Pred fuses U-Net and Transformer branches with decoders for motion, anisotropic Brownian components, and intensity. Diff-Gen uses ContextNet and conditional diffusion on the structure and residual, conditioned on the input and $\mathrm{Radar}^{\mathrm{pred}}$.
@@
Extended Data Figure 2: Nowcast3D’s layered forecasts of an extreme precipitation event in North China at 60, 120, and 180 min lead times, for altitudes ranging from 500 m to 13,000 m. This figure shows the layered prediction maps corresponding to Figure 2. We show a comparison between our prediction results and the ground truth (radar observations) at a specific vertical height. The observation at 500 m exhibits circular features due to the limitation of radar detection range, which appears as an overlap of multiple observation circles at low altitudes.
@@
Extended Data Figure 3: Nowcast3D’s layered forecasts of an extreme precipitation event in Maoming, Guangdong Province, China at 60, 120, and 180 min lead times, for altitudes ranging from 500 m to 13,000 m. This figure shows the layered prediction maps corresponding to in Figure 3. We show a comparison between our prediction results and the ground truth (radar observations) at a specific vertical height. The observation at 500 m exhibits circular features due to the limitations of the radar detection range, which appears as an overlap of multiple observation circles at low altitudes.
@@
Extended Data Figure 4: Wind field estimation of Nowcast3D forecast. a, Schematic of calculation from gridded data to site-specific wind field. b, Spatial averaging for evaluation: because observation sites do not exactly align with grid points, we define a circular area with a radius of Threshold around each site. The predicted values inside this circle are averaged for comparison with the site’s observation. c, Vertical averaging for evaluation: as observations may not match the exact vertical levels of the predictions, we average all observations within a vertical range of ±Threshold/2 at a specific height for comparison. d, Temporal evolution of the correlation coefficient between predicted and observed wind speeds. e, Temporal evolution of the mean absolute error in wind direction between predicted and observed winds.
Extended Data Figure 5: Radar network coverage and low-level imputation examples. a, Spatial distribution of the operational weather radar network over China and surrounding regions, with colored disks indicating approximate horizontal coverage at 500, 1500, 3000, 4500, 6500 and 10,000 m above the sea level. The dashed rectangle marks the analysis domain used in this study. b, Schematic of 3D radar sampling, illustrating how the beam widens with range and leads to sparse coverage near the surface despite dense sampling aloft. c, Example of low-level reflectivity in South China (the Huanan area) at attitudes of 500 m and 1,000 m before (left) and after (right) imputation. d, Same as c but for a convective event near Maoming, Guangdong Province, China, highlighting the recovery of coherent low-level structures beneath well-observed upper-level echoes.